OpenStudy (anonymous):

alright, need some help with this problem, not sure if im right or not: Evaluate the integral: ∫[(7π)/(6),(11π)/(6),4cscθcotθ,θ] calculating this, i got 0 as my answer... is that correct?

OpenStudy (anonymous):

\[\frac{\mathbb{d}}{\mathbb{d}x} \csc x = -\csc x \cot x \iff \int \csc x \cot x\ \mathbb{d}x = -\csc x \ +k\]

OpenStudy (anonymous):

Sorry, I didn't read your answer; yes, it equals 0.

OpenStudy (anonymous):

k, thanks~